The newspaper reported last week that Bennington Enterprises earned $34.18 million this year. The report also stated that the firm’s return on equity is 13 percent. Bennington retains 80 percent of its earnings.

  

What is the firm's earnings growth rate? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

  

  Earnings growth rate %

  

What will next year's earnings be? (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, e.g., 1,234,567. Round your answer to the nearest whole number, e.g., 32.)
  Next year’s earnings
Answer #1

Solution:

Return on equity (ROE) = 13% or 0.13

Retention ratio ( b) = 0.80

Current earning = $3,41,80,000

a)

Earnings growth rate = ROE *b

Earnings growth rate = 0.13*0.80= 0.104 or 10.40%

b) Next year earnings = Current earnings (1+ growth rate)

Next year earnings = $3,41,80,000*(1+0.104)= $3,77,34,720